OpenStudy (anonymous):

What is the x value for this system of equations? y=2-x 4x-3y=22 A.3 B.1 C.4 D.2

OpenStudy (mitu12):

it would be a because 4x-3(2-x)=22 will give you 3 for x

OpenStudy (anonymous):

substitute the fisrt equation in the second equation: 4x-3(2-x) = 22 4x - 6 + 3x = 22 7x = 22 + 6 7x = 28 x = 4 now plug it back on the first equation y = 2 - 4 y = -2 solution: (4, -2)

OpenStudy (anonymous):

option C

OpenStudy (mitu12):

i meant 4

OpenStudy (anonymous):

Oh, thank you both so much :)

OpenStudy (anonymous):

:)

OpenStudy (mitu12):

no problem